Two pumps are used to fill a water storage tank at a resort. One pump can fill the tank by itself in 9 hours, and the other can fill it in 6 hours. How long will it take both pumps operating together to fill the tank?
Answer by ptaylor(2198)

Let x=amount of time it takes both pumps to fill the pool when working together
So both pumps fill at the rate of 1/x tank per hour
First pump fills at the rate of 1/9 tank per hour
Second tank fills at the rate of 1/6 tank per hour
Together, they fill at the rate of 1/9 +1/6 tank per hour
So, our equation to solve is:
1/9 + 1/6=1/x multiply each term by 18x
2x+3x=18
5x=18
x=3 3/5 hours or 3h 36m--------------time it takes both tanks working together
CK
In 3 3/5 hours, tank one fills (1/9)(18/5)=2/5 of the tank
In 3 3/5 hours second tank fills (1/6)(18/5)=3/5 of the tank
2/5 + 3/5 =5/5 or the whole tank
Hope this helps---ptaylor